Is there a relationship between confidence intervals and two-tailed hypothesis tests? Let c be the level of confidence used to construct a confidence interval from sample data. Let α be the level of significance for a two-tailed hypothesis test. The following statement applies to hypothesis tests of the mean. For a two-tailed hypothesis test with level of significance α and null hypothesis H0: μ = k, we reject H0 whenever k falls outside the c = 1 − α confidence interval for μ based on the sample data. When k falls within the c = 1 − α confidence interval, we do not reject H0. (A corresponding relationship between confidence intervals and two-tailed hypothesis tests also is valid for other parameters, such as p, μ1 − μ2, or p1 − p2, which we will study later.) Whenever the value of k given in the null hypothesis falls outside the c = 1 − α confidence interval for the parameter, we reject H0. For example, consider a two-tailed hypothesis test with α = 0.01 and H0: μ = 20 H1: μ ≠ 20 A random sample of size 30 has a sample mean x = 23 from a population with standard deviation σ = 6. (a) What is the value of c = 1 − α? Construct a 1 − α confidence interval for μ from the sample data. (Round your answers to two decimal places.) lower limit Using methods of this chapter, find the P-value for the hypothesis test. (Round your answer to four decimal places.)
Answer #1

SOLUTION:

A)

We are given that =23 and  σ = 6, n=30

Level of significance α = 0.01.

H0: μ = 20

H1: μ ≠ 20

c = 1 − α=1-0.01=0.99

=2.58(6/)

=2.82

The end points are given by

-E=23-2.82=20.18

+E=25.82

We are 99% certain that the interval from 20.18 to 25.82 is an interval that contains the population mean time .

The value of given in the null hypothesis is 20.This value is not in the confidence interval we reject H0,, since the value fails outside of the specified range of the confidence interval .

B)

Level of significance α = 0.01

H0: μ = 20

H1: μ ≠ 20

Since,H1: μ ≠ 20 this is a two tailed test.Since the x distribution is normal and σ is known,use the standard normal distribution with

The sample of 30 measurements has mean =23,Converting this measurement to Z we have

=3/10.954=2.738

P-value=2P(Z<2.738)=2(0.003)=0.006

Since P-value 0.006<0.01,we see that P-value<α.

We reject H0.

The result are same in both the cases .In both the cases we reject the null hypothesis.